An aerosol can is empty, except for the propellant gas which exerts a pressure of 1.20 atm at 22 degrees Celsius. If the can is thrown into a fire and heated to a temperature of 535 degrees Celsius, what will be the pressure inside the hot can?


P= atm
Answer #1

we know, pilti - Pltz . Here Pi = 12 atm. sredot fo cus T1 = 22C = 22+273) X = 295K T ?!1015357.6 E 835+273)X = 808K izam x